CONFORMAL CHARACTERS AND THETA SERIES

Wolgang Eholzer Nils-Peter Skoruppa

Abstract. We describe the construction of vector valued modular forms transforming under a given congruence representation of the modular group SL(Z) in terms of theta series. We apply this general setup to obtain closed and easily computable formulas for conformal characters of rational models of W -algebras.

Typeset by AMS-TEX